OpenStudy (anonymous):

Please help! how do you rationalize 4/7+3i??

OpenStudy (anonymous):

4 / (7 + 3i) = (4(7 - 3i)) / ((7 + 3i)(7 - 3i)) = (28 - 12i) / (7^2 + 21i - 21i - 9i^2) = (28 - 12i) / (49 - 9(-1)) = (28 - 12i) / (49 + 9) = (28 - 12i) / 58 = (14 - 6i) / 29

OpenStudy (anonymous):

You're a life saver, thanks. I have one more, could you help?

OpenStudy (anonymous):

um...sure

OpenStudy (anonymous):

thanks, okay simply and change to radical form- (x^2/5)^3/4

OpenStudy (anonymous):

I got x^6/500...my teacher said I was wrong :/

OpenStudy (anonymous):